# Noncooperative Equilibrium Selection via a Trading-based Auction

### v1

@@ 6.3 Evaluation method
We evaluated the performance of TACo across three key metrics: optimality gap, fairness, and convergence stpng.
@@
  • Convergence stpng: The convergence stpng metric measures the algorithm’s efficiency by counting the number of stpng required for termination.
@@ 6.4 Results
Figure 3: Experimental results comparing the performance of the TACo algorithm with baseline methods (Voting, Utilitarian, Egalitarian, and Random dictator) across three evaluation metrics: optimality gap, Gini index, and termination stpng. TACo demonstrates the smallest median values for the optimality gap and Gini index, highlighting its effectiveness in achieving near-optimal and fair outcomes.
@@ 6.4.2 Termination proof validation
In our experiments, TACo reached consensus in a median of 53 stpng, with a maximum of 380 stpng. Given an ADS-B information exchange rate of 1 Hz, this corresponds to a median negotiation time of 53 seconds and a maximum of 6 minutes and 20 seconds. While TACo is iterative by design, these results suggest that its convergence remains within operationally feasible limits.
@@ 6.5 TACo variations for faster convergence
Figure 5: Effect of the decrement factor on TACo performance. (a) The termination stpng increase as the decrement factor $\gamma$ approaches 1, with a maximum exceeding 1500 stpng. For $\gamma\leq 0.9$, there is no notable reduction in convergence stpng. (b) The Gini index and the optimality gap remain largely unaffected by changes in $\gamma$, demonstrating that TACo maintains fairness and optimality regardless of the decrement factor.
(a) Termination stpng
@@
Figure 6: Effect of interruptions on TACo performance. (a) Interrupting the algorithm earlier reduces termination stpng effectively. (b) However, earlier interruptions increase the optimality gap and Gini index. This demonstrates the trade-off between faster convergence and reduced fairness and optimality.
@@ 6.5.1 Effect of decrement factor on convergence
We investigated how varying the decrement factor $\gamma$ impacts both convergence stpng and performance. In this experiment, we tested $\gamma$ values ranging from 0.3 to 0.99. As shown in Figure 5(a), the highest number of stpng occurred in the $\gamma=0.99$ case, with a median of 57 and a maximum of 1680 stpng. However, for $\gamma$ values below 0.9, there was no significant improvement in the convergence rate, even for the $\gamma=0.3$ case.
A similar trend was observed in the Gini index and optimality gap, as shown in Figure 5(b). Just as there was no significant improvement in the convergence rate for smaller $\gamma$ values, there was also no notable impact on optimality or fairness. Thus, as long as $\gamma$ does not exceed 0.9, there is no significant impact on convergence stpng, optimality, or fairness.
@@ 6.5.2 Effect of algorithm interruption on convergence
Figure 6 shows the results of these interruptions, comparing termination stpng and performance across varying levels of interruption. The results demonstrated that interruptions significantly reduced convergence stpng, as shown in Figure 6(a). However, as expected, forced consensus increased both the optimality gap and Gini index, as seen in Figure 6(b).

### v2

@@ 6.3 Evaluation method
We evaluated the performance of TACo across three key metrics: optimality gap, fairness, and convergence steps.
@@
  • Convergence steps: The convergence steps metric measures the algorithm’s efficiency by counting the number of steps required for termination.
@@ 6.4 Results
Figure 3: Experimental results comparing the performance of the TACo algorithm with baseline methods (Voting, Utilitarian, Egalitarian, and Random dictator) across three evaluation metrics: optimality gap, Gini index, and termination steps. TACo demonstrates the smallest median values for the optimality gap and Gini index, highlighting its effectiveness in achieving near-optimal and fair outcomes.
@@ 6.4.2 Termination proof validation
In our experiments, TACo reached consensus in a median of 53 steps, with a maximum of 380 steps. Given an ADS-B information exchange rate of 1 Hz, this corresponds to a median negotiation time of 53 seconds and a maximum of 6 minutes and 20 seconds. While TACo is iterative by design, these results suggest that its convergence remains within operationally feasible limits.
@@ 6.5 TACo variations for faster convergence
Figure 5: Effect of the decrement factor on TACo performance. (a) The termination steps increase as the decrement factor $\gamma$ approaches 1, with a maximum exceeding 1500 steps. For $\gamma\leq 0.9$, there is no notable reduction in convergence steps. (b) The Gini index and the optimality gap remain largely unaffected by changes in $\gamma$, demonstrating that TACo maintains fairness and optimality regardless of the decrement factor.
(a) Termination steps
@@
Figure 6: Effect of interruptions on TACo performance. (a) Interrupting the algorithm earlier reduces termination steps effectively. (b) However, earlier interruptions increase the optimality gap and Gini index. This demonstrates the trade-off between faster convergence and reduced fairness and optimality.
@@ 6.5.1 Effect of decrement factor on convergence
We investigated how varying the decrement factor $\gamma$ impacts both convergence steps and performance. In this experiment, we tested $\gamma$ values ranging from 0.3 to 0.99. As shown in Figure 5(a), the highest number of steps occurred in the $\gamma=0.99$ case, with a median of 57 and a maximum of 1680 steps. However, for $\gamma$ values below 0.9, there was no significant improvement in the convergence rate, even for the $\gamma=0.3$ case.
A similar trend was observed in the Gini index and optimality gap, as shown in Figure 5(b). Just as there was no significant improvement in the convergence rate for smaller $\gamma$ values, there was also no notable impact on optimality or fairness. Thus, as long as $\gamma$ does not exceed 0.9, there is no significant impact on convergence steps, optimality, or fairness.
@@ 6.5.2 Effect of algorithm interruption on convergence
Figure 6 shows the results of these interruptions, comparing termination steps and performance across varying levels of interruption. The results demonstrated that interruptions significantly reduced convergence steps, as shown in Figure 6(a). However, as expected, forced consensus increased both the optimality gap and Gini index, as seen in Figure 6(b).